OpenStudy (anonymous):

Find 3 consecutive odd numbers where the product of the smaller two numbers is 34 less than the square of the largest number.

OpenStudy (anonymous):

you can say the consecutive number is x

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

so three odd number are x , x+2 and x+4

OpenStudy (anonymous):

now you can say the product of 2 smaller number is x(x+2)

OpenStudy (anonymous):

square of the largest number is (x+4)^2 and less than 34 is (x+4)^2-34

OpenStudy (anonymous):

so it will be x(x+2) = (x+4)^2-34

OpenStudy (anonymous):

can you solve now

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

x^2+2x = x^2 + 2(x)(4) + (4)^2 -34 x^2+2x = x^2 + 8x +16 -34 x^2-x^2 = 8x - 2x +16 -34 18 = 6x 18 / 6 = x x= 3